Find f(3) when f(x)=x^2-4x+7
Found 2 solutions by MathLover1, nerdybill:
Answer by MathLover1(20849)
Answer by nerdybill(7384)

Find f(3) when f(x)=x^2-4x+7
.
just substitute the x with a 3:
f(x)=x^2-4x+7
f(3)=3^2-4(3)+7
f(3)=9-4(3)+7
f(3)=9-12+7
f(3)=-3+7
f(3)=4